OpenStudy (anonymous):

Round all answers to the nearest hundredh. Evaluate each expresion for the values given.. b^2-4ac;b=3.a=-1,c=-2

OpenStudy (anonymous):

just put the values

OpenStudy (anonymous):

im so lost.. i been on this for twoo days

OpenStudy (anonymous):

3^2-4(-1)(-2)

OpenStudy (anonymous):

can you do it now

OpenStudy (anonymous):

10

OpenStudy (anonymous):

3^2 = 9 4(-1)(-2) = 8 now do it

OpenStudy (anonymous):

not 10 its 9-8 =???

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

b^2 - 4ac Replace the letters with their given values and do the math. (3)^2 - 4(-1)(-2) 9 - 8 = 1 The answer is 1. Look for me on facebook for picture reply to math questions.